OpenStudy (anonymous):

If the midpoint between (x, 7) and (10, 11) is (15, 9), find the value of x

jimthompson5910 (jim_thompson5910):

The x coordinate of the midpoint is the average of the two x coordinates of the endpoints This means that (x+15)/2 = 10 Solve that for x to get your answer

OpenStudy (anonymous):

i dont understand how to solve that..

jimthompson5910 (jim_thompson5910):

(x+15)/2 = 10 x+15 = 10*2 x+15 = 20 x = ???

OpenStudy (anonymous):

x= 5 :) thank you lol!

jimthompson5910 (jim_thompson5910):

you got it